OpenStudy (anonymous):

multiply: (9+2i)(8-5i)

OpenStudy (anonymous):

you are going to use the FOIL method, so we have 4 products we need to get: First: 9*8 = 72 Outter: 9*(-5i) = -45i Inner: (2i)*8 = 16i Last: (2i)*(-5i) = -10i^2 Now we just add everything up and get: 72 -45i+16i-10i^2 = 72 - 29i - 10i^2 but i^2 is -1, so now we have: 72 - 28i - 10(-1) = 72 - 29i + 10 = 82 - 29i

OpenStudy (anonymous):

Use Normal Multiplication And Value i^2 =-1